Amanda Rice has just arranged to purchase a $460,000 vacation home in the Bahamas with a 25 percent down payment. The mortgage has a 5.2 percent APR compounded monthly and calls for equal monthly payments over the next 30 years. Her first payment will be due one month from now. However, the mortgage has an eight-year balloon payment, meaning that the balance of the loan must be paid off at the end of Year 8. There were no other transaction costs or finance charges. How much will Amanda’s balloon payment be in eight years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Answer:

Loan amount = Cost of vacation home - down payment = $460,000 - 25% * 460000 = $345,000

Monthly interest rate = 5.2%/12

Number of monthly installments = 30 years * 12 = 360

First payment due on one month from now

To get monthly payments, we will use PMT function of excel:

PMT (rate, nper, pv, fv, type)

= PMT (5.2%/12, 360, -345000, 0, 0)

= $1,894.432539

The outstanding balance at the end of year 8 is equal to PV of remaining monthly payments

Balance number of monthly payments = 360 - 8 * 12 =264

Hence PV at the end of year 8:

PV (rate, nper, pmt, fv, type)

= PV (5.2%/12, 264, -1894.432539,0,0)

=$297,572.85

Amanda’s balloon payment be in eight years = $297,572.85